OpenStudy (anonymous):

A fan at a rock concert is 50.0 m from the stage, and at this point the sound intensity level is 108dB . Sound is detected when a sound wave causes the tympanic membrane (the eardrum) to vibrate . Typically, the diameter of this membrane is about 8.40 mm in humans How much energy is transferred to her eardrums each second ? How fast would a 2.00 mg mosquito have to fly to have this much kinetic energy? How fast a typical 2.0 mg mosquito would have to fly (in mm/s) to have an amount of energy delivered to the eardrum each second when someone whispers (20 dB) a secret in your ear

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

Help please!!!

OpenStudy (anonymous):

Help anyone?

OpenStudy (anonymous):

Do you remember equation for sound intensity?

OpenStudy (anonymous):

Is it 10log(I/10^-12)?

OpenStudy (anonymous):

Yes, we call that intensity level (known in the problem). Can you find I (intensity) from than eq?

OpenStudy (anonymous):

I got .0631?

OpenStudy (anonymous):

Explain your intermediate step?

OpenStudy (anonymous):

(108/10)=10.8 10^10.8=6.31 6.31*10^-12=.0631=I

OpenStudy (anonymous):

Where did you get 6.31?

OpenStudy (anonymous):

To get rid of log, don't you you raise 10 by 10.8?

OpenStudy (anonymous):

Yep, correct. Just my habit to make sure everything goes right :) Now you know I, remember what I is?

OpenStudy (anonymous):

.0631

OpenStudy (anonymous):

I mean the physical definition or the equation for I

OpenStudy (anonymous):

Power/Area

OpenStudy (anonymous):

Great, can you find the power of sound that mr rock fan receive?

OpenStudy (anonymous):

how do i know the area though?

OpenStudy (anonymous):

the area here means the area of sphere located 50 m from the source. Or simply area of sphere with radii 50 m.

OpenStudy (anonymous):

I got 31415.93 for the area? If you could stay here, I'd appreciate it. I'll be back in 10-15 minutes.

OpenStudy (anonymous):

Yep, now find the power.

OpenStudy (anonymous):

1982.21

OpenStudy (anonymous):

Find the energy ear receive from this power. You know what power is right?

OpenStudy (anonymous):

I have no clue how to find the energy ear from power

OpenStudy (anonymous):

you already know P. What is power?

OpenStudy (anonymous):

the rate that energy is transferred. not sure on what the equation is

OpenStudy (anonymous):

The first question asked is energy ear receive each second. Or the power received by ear.

OpenStudy (anonymous):

So, do I have to find the area of the ear and multiply that by I and find that power?

OpenStudy (anonymous):

To be clear, the question asked for intensity again, but now only the amount received by ear. Think like this: the 'sphere' has known power. What is part of power received by some tiiiny area (eardrum's area) of that spehere?

OpenStudy (anonymous):

Sorry, but I just don't understand. If you're saying find the power of the ear, I got 1.39*10^-5

OpenStudy (anonymous):

It's just the previous I times ear area.

OpenStudy (anonymous):

ok, then I got that above your post

OpenStudy (anonymous):

Then to mosquito?

OpenStudy (anonymous):

I got the answer wrong. The answer is supposed to be in microjoules so I tried converting and ended up with 1.39*10^-11 and it isn't right. Did I do something wrong? I only have two tries left to get this particular question right

OpenStudy (anonymous):

I found 3,5 uJ

OpenStudy (anonymous):

(for energy transferred)

OpenStudy (anonymous):

is that 3.5? the comma threw me off

OpenStudy (anonymous):

Sorry, yes, 3.5. in my country, we use 3,5 instead.

OpenStudy (anonymous):

Thank You! Could you help me with the mosquito one next?

OpenStudy (anonymous):

You are in the right track, just miscalculated i guess. For the mosquito, remember the equation for kinetic energy? If you do, the kinetic energy must be equal to 3,5 uJ.

OpenStudy (anonymous):

in uJ or just Joules?

OpenStudy (anonymous):

Whichever you want, just dont forget to equate the unit.

OpenStudy (anonymous):

Can anyone help me with this second question?

OpenStudy (anonymous):

kinetic energy = 1/2mv^2 correct? Because I'm not sure if i'm getting the right answer

OpenStudy (anonymous):

K=1/2 m v^2 right. This K must be equal to your 3,5 uJ :)

OpenStudy (anonymous):

ok, got it. so the third question?

OpenStudy (anonymous):

Just redo the whole problew with that new intensity.

OpenStudy (anonymous):

I got 886.23 but got it wrong?

OpenStudy (anonymous):

For the power received by ear, just multiply 20dB with ear area. Make sure the units are correct. Do this once again, i'm sorry. I have to sleep now, already 2 am here.

OpenStudy (anonymous):

Ok, I keep getting the third question wrong. Can someone help me?

OpenStudy (anonymous):

What's your last answer for third question.

OpenStudy (anonymous):

I had .1489

OpenStudy (anonymous):

What is the intensity with 20dB intensity level?

OpenStudy (anonymous):

1*10^-10

OpenStudy (anonymous):

Unit?

OpenStudy (anonymous):

W/m^2

OpenStudy (anonymous):

Okay, now multiply this with ear area in m^2

OpenStudy (anonymous):

I got 2.217*10^-14, is that correct so far?

OpenStudy (anonymous):

I found 5,541 x 10^-15. Area of eardrum in m^2=?

OpenStudy (anonymous):

For the area of the eardrum I got 4pi(.0042^2)=2.2167*10^-4. Is my formula right?

OpenStudy (anonymous):

I got 5,542 x 10^-5 m^2 for area. "4pi"?

OpenStudy (anonymous):

Are we calculating the area of a sphere? or is it supposed to be a circle?

OpenStudy (anonymous):

It's a circle I guess.

OpenStudy (anonymous):

Oh, okay then. So, then I would calculate power and end up with 5.542*10^-15

OpenStudy (anonymous):

Yes, that's energy in one scond. The unit's still Joule, to keep our track of unit in SI. Now find the speed of mosquito

OpenStudy (anonymous):

I got 7.44*10^-5 m/s. How would I get this to mm/s?

OpenStudy (anonymous):

You remember to use kg for mass?

OpenStudy (anonymous):

yes, I did that. 2*10^-6 for the mosquito's mass

OpenStudy (anonymous):

Okay, to convert to mm/s, just multiply with 10^3

OpenStudy (anonymous):

Thanks! I have a fourth question for this problem and it says "Compare the mosquito's speeds found in parts B and C" and then it says "Please enter your number (factor) without the units." I have no idea what this is asking.

OpenStudy (anonymous):

it means vb/vc Dont forget to make the units same first.

OpenStudy (anonymous):

Thanks so much for your help. Would it be ok with you if I asked a couple more open questions. I have one chance left to get those right and I feel like I did them right

OpenStudy (anonymous):

Yes, i'd like to help. But i have like 15 minutes. Have to do something else :)